OpenStudy (imterribleatmath):

Let f(x)=3-4x..... find x when f(x)=-2... I know the answer is -1.25, or I hope it is.. But how do you get that answer?

OpenStudy (ajprincess):

Equate both equations. 3-4x=-2 See if u can do nw @imterribleatmath

OpenStudy (imterribleatmath):

What do you mean by equate? Subtract the 3 on both sides?

OpenStudy (imterribleatmath):

@ajprincess

OpenStudy (ajprincess):

f(x)=3-4x f(x)=-2 Equate the equations. then u get 3-4x=-2 Solve for x nw. Getting it?

OpenStudy (anonymous):

when lim x intend to 0 then proove that log cosx/ sin2x =1/2

OpenStudy (ajprincess):

@imterribleatmath

OpenStudy (anonymous):

can u help mee to solve the problem

OpenStudy (ajprincess):

@tirthdutta21 plz post ur question as a new one. I will try my best to help u.:)

OpenStudy (ajprincess):

Well do u knw how to post a new question @tirthdutta21?

OpenStudy (imterribleatmath):

I'm trying to solve for x... 3-4x=-2 subtract by -3 on both sides? -4x=-5 divide by -4? -5/-4... It gives me 1.25.. The answer should be -1.25 right? @ajprincess thanks for your help btw

OpenStudy (ajprincess):

if ur question is 3-4x=-2 then x=1.25 as far as I knw

OpenStudy (ajprincess):

nd welcome:)

OpenStudy (imterribleatmath):

a yes, I was using my calculator wrong.. oops lol

OpenStudy (ajprincess):

oh that's k:)